# Homological data on the periodic structure of self-maps on wedge sums

### v1

@@ 4. Self-maps on wedge sum of tori
Since the Lefschetz zeta functions are rational functions, we define its degree as the degree of the numerator minus the degree of the denominator. For toral maps with non zero eigenvalues, the degree of their Lefschetz zeta function is zero [5]. Notice that in this case the maps $f_{ii}$, are toral maps defined on $\mathbb{T}^{n}$. Hence the degree of
@@
is zero. Therefore the degree of $\zeta_{f}(t)$ is $-1$, which implies that $\zeta_{f}(t)$ is not constant, i.e. the map $f$ is not Lefschetz periodic point free. ∎
@@
Let $n$ be an odd prime Let $g_{c}:\mathbb{T}^{n}\to\mathbb{T}^{n}$ be a continuous map such that the characteristic polynomial of the induced map $(g_{c})_{*1}$ is $t^{n}-c$, with $c>2$ a positive integer. Let $f:\bigvee_{i=1}^{s}\mathbb{T}^{n}\to\bigvee_{i=1}^{s}\mathbb{T}^{n}$, a permutative map, with permutation $\sigma$, such that its coordinates are given by: $f_{j\sigma(j)}=g_{c_{j}}$, and $f_{jk}$ is the constant map equal to the base point for $k\neq\sigma(j)$, and $c_{j}>2$ a positive integer, for $1\leq q\leq s$. Then $\mbox{APer}_{L}(f)=\mathbb{N}$.
@@
In [7], it was proved that $\ell(g_{c}^{m})<0$, for all $m>0$, when $c>2$ and $n$ an odd prime. From Theorem 2 follows that the set of the algebraic of $f$ is the set of the positive integers. ∎
In [7], it was conjectured that $\ell(g_{c}^{m})<0$, for all $m>0$, when $c>2$ any positive integer $n>1$. Therefore we conjecture that the Corollary 12 is true for all positive $n$.
@@ 5. Examples
Example 1: Consider the self-map $f$ on $\mathbb{T}^{2}\vee\mathbb{T}^{2}$, such that the coordinate-maps $f_{11}$, $f_{1,2}$, $f_{2,1}$, $f_{2,2}$ are given by the toral maps whose matrix representation are
@@
$L(f^{m})=1-\text{trace}(f_{*1}^{m})+\text{trace}(f_{*2}^{m})=\left\{\begin{array}[]{ll}-1,&m=0\pmod{4}\\ 1,&m=1,3\pmod{4}\\ 7,&m=2\pmod{4}.\end{array}\right.$
@@
If can be easily checked that algebraic periods are all the even numbers.
Example 3: Consider the self-map $f$ on $\mathbb{T}^{2}\vee\mathbb{T}^{2}$, such that the coordinates $f_{11},\,f_{1,2},\,f_{2,1},\,f_{2,2}$ are given by the toral maps with matrix representation:
@@
$f_{*2}:\left\{\begin{array}[]{ll}\gamma_{1}^{1}\wedge\gamma_{1}^{2}\to f_{*1}(\gamma_{1}^{1})\wedge f_{*1}(\gamma_{1}^{2})&=-\gamma_{1}^{2}\wedge(\gamma_{2}^{2}-\gamma_{1}^{2})=-\gamma_{1}^{2}\wedge\gamma_{2}^{2}+\gamma_{1}^{2}\wedge\gamma_{1}^{2}=-\gamma_{1}^{2}\wedge\gamma_{2}^{2}.\\ \gamma_{2}^{1}\wedge\gamma_{2}^{2}\to f_{*1}(\gamma_{2}^{1})\wedge f_{*1}(\gamma_{2}^{2})&=(\gamma_{1}^{1}-\gamma_{1}^{2})\wedge(\gamma_{2}^{1}-\gamma_{1}^{2})=\\ &=\gamma_{1}^{1}\wedge\gamma_{2}^{1}-\gamma_{1}^{1}\wedge\gamma_{1}^{2}-\gamma_{1}^{2}\wedge\gamma_{2}^{1}+\gamma_{1}^{2}\wedge\gamma_{1}^{2}=\\ &=\gamma_{1}^{1}\wedge\gamma_{2}^{1}-\gamma_{1}^{1}\wedge\gamma_{1}^{2}-\gamma_{1}^{2}\wedge\gamma_{2}^{1}.\\ \end{array}\right.$
@@
Examples 3 and 4 show that not any $4\times 4$ integer matrix gives rise to a map $f:\mathbb{T}^{2}\vee\mathbb{T}^{2}\to\mathbb{T}^{2}\vee\mathbb{T}^{2}$ having this matrix as $f_{*1}$.
@@
$L(f^{m})=1-\text{trace}(f_{*1}^{m})+\text{trace}(f_{*2}^{m})=\left\{\begin{array}[]{ll}0,&m\equiv 0\pmod{4}\\ 2,&m\equiv 1,3\pmod{4}\\ 4,&m\equiv 2\pmod{4};\end{array}\right.$
@@
Its Dold coefficients are $\ell(f)=\ell(f^{2})=2$, $\ell(f^{4})=-2$ and $\ell(f^{m})=0$ if $m\neq 1,2,4$. Hence the algebraic periods set is $\mbox{APer}_{L}(f)=\{1,2,4\}.$
@@ 6. Open questions and further remarks
    Examples 3 and 4 show that not any $4\times 4$ integer matrix allows to define a map $f:\mathbb{T}^{2}\vee\mathbb{T}^{2}\to\mathbb{T}^{2}\vee\mathbb{T}^{2}$ having this matrix as $f_{*1}$, more generally not any coordinate maps give a well-defined map on the wedge sum. This give rises to the following questions:
@@
      Let $(X,x_{0})$ be a pointed topological manifold, and $s\times s$ continuous maps $f_{ij}:X\to X$. Under which conditions the maps $f_{ij}$ can produce a continuous self-map $f$ on $\bigvee_{j=1}^{s}X$ having the maps $f_{ij}$ as coordinates?
      Example 5 shows that there are continuous and non-trivial self-map self-maps on $\bigvee_{i=1}^{s}\mathbb{T}^{n}$, which are not of permutation type.
@@
      In the particular case of $X=\mathbb{T}^{n}$: Which matrices of integer coefficients arise from self-maps of the wedge sum of the tori? We point out, that for graph maps, i.e. $n=1$, there are not such obstructions, any possible integer matrix of the right dimension is allowed, since all the homology groups of $H_{k}(\bigvee_{i=1}^{s}{\mathbb{S}}^{1},\mathbb{Q})$ are trivial for $k>1$.
@@
    The main difference between the study of the dynamics of the maps on the wedge sum of tori, of dimension larger than $1$ with the graphs maps, is that for the later there are no obstructions on the homology groups of dimension bigger than $1$, which gives no obstructions on the possible maps on homology, as in the toral case, that was described in the point (1).
@@
    Give general conditions for a map $f:\bigvee_{i=1}^{s}X_{i}\to\bigvee_{i=1}^{s}X_{i}$ to be a Lefschetz periodic point free and partially periodic point free. Partial answers to this question have been given depending the nature of the spaces $X_{i}$. In the case of topological graphs and wedge sum of spheres it was studied partially before, for instance see [26, 29, 33].
@@
    Give more precise description of the set of algebraic periods of permutative toral maps.

### v2

@@ 4. Self-maps on wedge sums of tori
Since Lefschetz zeta functions are rational functions, we define its degree as the degree of the numerator minus the degree of the denominator. For toral maps with non zero eigenvalues, the degree of their Lefschetz zeta function is zero [5]. Notice that in this case the maps $f_{ii}$, are toral maps defined on $\mathbb{T}^{n}$. Hence the degree of
@@
is zero. Therefore the degree of $\zeta_{f}(t)$ is 1-1, which implies that $\zeta_{f}(t)$ is not constant, i.e. the map $f$ is not Lefschetz periodic point free. ∎
@@
Let $n$ be an odd prime. Let $g_{c}:\mathbb{T}^{n}\to\mathbb{T}^{n}$ be a continuous map such that the characteristic polynomial of the induced map $(g_{c})_{*1}$ is $t^{n}-c$, with $c>2$ a positive integer. Let $f:\bigvee_{i=1}^{s}\mathbb{T}^{n}\to\bigvee_{i=1}^{s}\mathbb{T}^{n}$ be a permutative map with permutation $\sigma$ such that its coordinates are given by: $f_{j\sigma(j)}=g_{c_{j}}$ and $f_{jk}$ is the constant map equal to the base point for $k\neq\sigma(j)$, and $c_{j}>2$ a positive integer, for $1\leq q\leq s$. Then $\mbox{APer}_{L}(f)=\mathbb{N}$.
@@
In [7], it was proved that $\ell(g_{c}^{m})<0$, for all $m>0$, when $c>2$ and $n$ an odd prime. From Theorem 2 follows that the set of the algebraic integers of $f$ is the set of the positive integers. ∎
In [7], it was conjectured that $\ell(g_{c}^{m})<0$, for all $m>0$, when $c>2$ any positive integer $n>1$. Therefore we conjecture that Corollary 12 holds true for all positive $n$.
@@ 5. Examples
Example 1: Consider the self-map $f$ on $\mathbb{T}^{2}\vee\mathbb{T}^{2}$, such that the coordinate-maps $f_{1,1}$, $f_{1,2}$, $f_{2,1}$, $f_{2,2}$ are given by the toral maps whose matrix representation are
@@
[…]
@@
It can be easily checked that algebraic periods are all the even numbers.
Example 3: Consider the self-map $f$ on $\mathbb{T}^{2}\vee\mathbb{T}^{2}$, such that the coordinates $f_{1,1},\,f_{1,2},\,f_{2,1},\,f_{2,2}$ are given by the toral maps with matrix representation:
@@
[…]
@@
Examples 3 and 4 show that not any $4\times 4$ integer matrix gives rise to a map $f:\mathbb{T}^{2}\vee\mathbb{T}^{2}\to\mathbb{T}^{2}\vee\mathbb{T}^{2}$ having such matrix as $f_{*1}$.
@@
[…]
@@
Its Dold coefficients are $\ell(f)=\ell(f^{2})=2$, $\ell(f^{4})=-2$, and $\ell(f^{m})=0$ if $m\neq 1,2,4$. Hence the set of algebraic periods is $\mbox{APer}_{L}(f)=\{1,2,4\}.$
@@ 6. Open questions and further remarks
    Examples 3 and 4 show that not any $4\times 4$ integer matrix allows to define a map $f:\mathbb{T}^{2}\vee\mathbb{T}^{2}\to\mathbb{T}^{2}\vee\mathbb{T}^{2}$ having this matrix as $f_{*1}$, more generally not any coordinate maps give a well-defined map on a wedge sum. This gives rise to the following questions:
@@
      Let $(X,x_{0})$ be a pointed topological manifold, and $s\times s$ continuous maps $f_{ij}:X\to X$. Under which conditions the maps $f_{ij}$ produce a continuous self-map $f$ on $\bigvee_{j=1}^{s}X$ having the maps $f_{ij}$ as coordinates?
      Example 5 shows that there are continuous and non-trivial self-maps on $\bigvee_{i=1}^{s}\mathbb{T}^{n}$, which are not of permutation type.
@@
      In the particular case of $X=\mathbb{T}^{n}$: Which matrices of integer coefficients arise from self-maps on wedge sums of the tori? We point out, that for graph maps, i.e. $n=1$, there are no such obstructions, any possible integer matrix of the right dimension is allowed, since all the homology groups of $H_{k}(\bigvee_{i=1}^{s}{\mathbb{S}}^{1},\mathbb{Q})$ are trivial for $k>1$.
@@
    The main difference between the study of the dynamics of maps on wedge sums of tori, of dimension larger than $1$ with graphs maps, is that for the latter there are no obstructions on the homology groups of dimension bigger than $1$.
@@
    Give general conditions for a map $f:\bigvee_{i=1}^{s}X_{i}\to\bigvee_{i=1}^{s}X_{i}$ to be Lefschetz periodic point free and partially periodic point free. Partial answers to this question have being given depending the nature of the spaces $X_{i}$. In the case of topological graphs and wedge sum of spheres it was studied partially before, for instance see [26, 29, 33].
@@
    Give a more precise description of the set of algebraic periods of permutative toral maps.
Acknowledgements: The authors would like to thank the anonymous referees for their remarks which helped to improve the article.